a right triangle is such that the shortest side is 7 inches shorter than the longest and the other side is only 1 inch shorter than the longest side.find the length of the 3 sides
Answer by Boreal(15235)

x is shortest side
x+7 is longest
x+6 is the third
x^2+(x+6)^2=(x+7)^2
x^2+x^2+12x+36=x^2+14x+49
x^2-2x-13=0
x=(1/2)(2+/- sqrt (4+52)
-
x=1+sqrt(14)
x+6=7+sqrt(14)
hypot=8+sqrt(14)
1+2 sqrt(14)+14
49+14sqrt(14)+14=
64+16 sqrt(14)+14, as it should be.
-
side lengths are 1+ sqrt(14), 7 +sqrt (14), and 8 +sqrt (14)